40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 2353     Entry time: Mon Jun 3 17:44:30 2019
Author: anchal 
Type: DailyProgress 
Category: NoiseBudget 
Subject: New Beat Note measurement 

Today I took a beat note noise measurement to see where we are. Attached is the updated noise budget.


Measurement Setup:

  • I used wideband New Focus 1811 for detecting beat note. I was using the resonant Beat Note Detectot SN 101 only. I must have misinterpreted the connection as I was seeing about -1 dBm signal even after the BN detector being resonant. And that must have happened because we have increased the light intensity by a lot.
  • The cavity transmitted power levels are 5.2 mW (South) and 4.7 mW (North) with ~70% and ~60% mode matching respectively.
  • I didn't play around much with FSS gains but as of now, COM gain values are -0.3V (0.4 dB) and 2.5V (90 dB) in North and South Paths respectively. Fast Gain values are 1.4 V (54.8 dB) and 1.15 V (46.8 dB) respectively.
  • Marconi Actuation slope was set to 10KHz/V with carrier frequency drifting at less than 20 kHz/min. The carrier frequency was near 300 MHz, so it was outside the bandwidth of any detector we have. This measurement was completely bogus!
  • Gain of SR560 in PLL was set to 200.
  • With this, the bandwidth of the PLL was measured as 76.5kHz.
  • I took 4 different FFT sweeps at different ranges with averaging such that the measurement remains between carrier frequency jump. For now, I was doing the trigger manually by hand, but we intend to make this automated

Analysis and Result:

  • I'm still using the old noise budget notebook for plotting the results because of the unresolved issues mentioned in CTN:2277. I'll work on this as soon as we get to the point where careful analysis of this noise budget is indeed important.
  • Clearly, we aren't doing any better. I'm thinking of a new frequency noise measurement scheme which I discussed with Andrew before he left. This scheme might help us let go of any temperature control whatsoever.
  • More importantly, though, I definitely need to fix whatever is kicking the FSS and not letting me increase the gain of the feedback loop.

Edit Thu Jun 6 14:17:34 2019 Anchal : Above in red.


Edit Tue Aug 13 17:08:46 2019 anchal:

The PLL Readout noise added on this plot was erroneous and I can't find where it came from either. So the noisebudget attached is wrong! I was a dumbo then.

Attachment 1: 20190603_174101noiseBudget.pdf  943 kB  Uploaded Mon Jun 3 19:03:45 2019  | Hide | Hide all
20190603_174101noiseBudget.pdf
ELOG V3.1.3-